# A Survey on Web Application Testing: A Decade of Evolution

### v1

@@ 1. Introduction
The rest of this paper is organized as follows: Section 2 provides background information on the formal definition of WAT and the fundamental testing techniques employed in the field. Section 3 outlines the survey methodology, including presenting the six research questions, the literature retrieval process, and the statistical analysis of the retrieved studies. Sections 4 to 9 address each of the research questions (RQ1 to RQ6), respectively. Finally, Section 10 concludes the paper by summarizing key findings and proposing directions for future research.
@@ 5.2.3. Performance Evaluation
Advanced tools such as JMeter (Kiran et al., 2015; Tiwari et al., 2023) and LoadRunner (Cai and Hu, 2014) can automate measurement of some important metrics, including response times, throughput, and system resource utilization (e.g., CPU, memory, disk I/O) (Fan et al., 2023). They can do this across a spectrum of load levels (Fan et al., 2023). These tools can provide comprehensive insights into system behavior under stress, helping the optimization of the system architecture, and ensuring reliability during peak load conditions (Leotta et al., 2020).
@@ 7.3.1. Application of Effectiveness and Efficiency Metrics in WAT
Efficiency metrics are not only used for coverage and fault-detection capabilities, but also in execution time and resource utilization (Putri et al., 2017; Ali and Badr, 2015). Empirical studies have shown that tools like JMeter can optimize performance testing through distributed architectures (Saharan et al., 2018; Kiran et al., 2015). By parallelizing the execution of test cases, JMeter can significantly reduce execution time while ensuring system stability under high-load conditions. This architecture has been shown to perform well in real-world scenarios, especially in large-scale multi-user concurrent environments (Guarnieri et al., 2017).

### v2

@@ 1. Introduction
The rest of this paper is organized as follows: Section 2 provides background information on the formal definition of WAT and the fundamental testing techniques employed in the field. Section 3 outlines the survey methodology, including presenting the six research questions, the literature retrieval process, and the statistical analysis of the retrieved studies. Sections 4 to LABEL:SEC:rq8 address each of the research questions (RQ1 to RQ6), respectively. Finally, Section 10 concludes the paper by summarizing key findings and proposing directions for future research.
@@ 5.2.3. Performance Evaluation
Advanced tools such as JMeter (Kiran et al., [n. d.]; Tiwari et al., 2023) and LoadRunner (Cai and Hu, 2014) can automate measurement of some important metrics, including response times, throughput, and system resource utilization (e.g., CPU, memory, disk I/O) (Fan et al., 2023). They can do this across a spectrum of load levels (Fan et al., 2023). These tools can provide comprehensive insights into system behavior under stress, helping the optimization of the system architecture, and ensuring reliability during peak load conditions (Leotta et al., 2020).
@@ 7.3.1. Application of Effectiveness and Efficiency Metrics in WAT
Efficiency metrics are not only used for coverage and fault-detection capabilities, but also in execution time and resource utilization (Putri et al., 2017; Ali and Badr, 2015). Empirical studies have shown that tools like JMeter can optimize performance testing through distributed architectures (Saharan et al., 2018; Kiran et al., [n. d.]). By parallelizing the execution of test cases, JMeter can significantly reduce execution time while ensuring system stability under high-load conditions. This architecture has been shown to perform well in real-world scenarios, especially in large-scale multi-user concurrent environments (Guarnieri et al., 2017).